# Normalized solutions of Nehari-Pankov type to indefinite variational problems

### v1

@@ 1. Introduction
$Au=\lambda u+I^{\prime}(u),\qquad u\in E$ (1)
@@
$\|u\|_{H}=\mu$ (2)
for some given $\mu>0$. This constrained problem also has a variational formulation: its solutions are precisely the critical points of the restriction of the energy functional $\mathcal{J}_{0}$ to the sphere $S_{\mu}:=\{u\in H\>:\>\|u\|_{H}=\mu\}$, and the parameter $\lambda$ appears as an a priori unknown Lagrange multiplier.
@@
$u^{\prime\prime}=\lambda u+f(u)$ (3)
@@
    $F\Bigl(\frac{s}{\kappa_{0}}\Bigr)\leq F(-s)\leq F(\kappa_{0}s)\qquad\text{for all $s\geq s_{0}$.}$ (4)
@@
Very recently, in [16, Theorem 1.3], Jeanjean and Song proved a multiplicity result for compact graphs $\mathcal{G}$ and small masses with continuity-Kirchhoff conditions. More precisely, by combining a localized variant of Lusternik-Schnirelman theory based on flow invariance with delicate estimates, they show that for every $n\in\mathbb{Z}^{>0}$ there exists $\mu_{n}>0$ with the property that, for every $\mu\in(0,\mu_{n})$, the NLS equation $u^{\prime\prime}=\lambda u+|u|^{p-2}u$ admits at least $n$ pairs $\pm u$ of sign changing solutions with fixed mass $\mu$. Using our new $\lambda$ to masses” tool (Theorem 1.7 below), we are able to improve this multiplicity result to the existence of infinitely many solutions, for any mass. Moreover, we can extend this multiplicity result beyond the class of odd nonlinearities to a setting where Lusternik-Schnirelman theory does not apply. Let us also point out that we can deal with the most general self-adjoint vertex conditions, also generalizing the previous results. In the case of continuity-Kirchhoff conditions, the first eigenfunction of $A$ is positive in $\mathcal{G}$ (see e.g. [18, Theorem 4.12]), so that we also obtain the existence of sign-changing solutions. Finally, as already pointed out, we do not need to make any case distinction depending on whether $p<6$ or $6<p$.
@@
$(-\Delta)^{m}u=\lambda u+f(x,u)\qquad\text{in }\Omega$ (5)
@@
$u=\partial_{\nu}u=\dots=\partial_{\nu}^{m-1}u=0\qquad\text{on }\partial\Omega$ (6)
@@
$u=\Delta u=\cdots=\Delta^{m-1}u=0\qquad\text{on }\partial\Omega.$ (7)
@@
Suppose that (FP1)–(FP4) are satisfied. Then, for any integer $k\geq 1$, there exists $\mu_{k}>0$ such that for all $\mu\in\mathopen{(}0,\mu_{k}\mathclose{)}$, problems (5)–(6) and (5)–(7) both admit $k$ distinct pairs of solutions $(\pm u,\lambda)$ with mass $\|u\|_{L^{2}(\Omega)}^{2}=\mu$. Moreover, if $m=1$, these solutions are sign-changing.
We emphasize that this result is new even in the classical second order case $m=1$, and it is the precise analogue of the recent multiplicity result of Jeanjean and Song in [16] for the case of compact metric graphs and the nonlinearity $f(t)=|t|^{p-2}t$. We point out that our method is completely different from the approach in [16], which, according to remarks in [16, Introduction], also yields multiplicity results for NLS-equations on bounded domains in $\mathbb{R}^{N}$.
We also note that the conclusion of Theorem 1.3 is weaker than the one of Theorem 1.1, since only small masses are considered. We conjecture that at least in the case $N=2$, $m>1$ the multiplicity result of Theorem 1.3 extends to arbitrary positive masses $\mu$. To support this conjecture, we present a further result for the closely related problem of finding normalized solutions of biharmonic semilinear equations in the $2$-torus.
@@
$\Delta^{2}u=\lambda u+r(x)|u|^{p-2}u\qquad\text{in $\Omega$}$ (8)
@@
In the mass supercritical regime $p>6$ and for a given positive mass $\mu>0$, even the mere existence of one normalized solution of this problem is new. If the weight function $r$ is constant, then the symmetries of the torus can be used to produce solutions with large mass via a reflection procedure (see [21, Theorem 1.13 and Remark 1.1] for related results in bounded domains). This is not possible when $r$ does not have any symmetries.
@@
    $\lim_{t\to 0^{+}}\frac{I(tu)}{t^{2}}=0\qquad\text{and}\qquad\lim_{t\to+\infty}\frac{I(tu)}{t^{2}}=+\infty.$ (9)
@@
    $I(u)-I((1+s)u+v)+I^{\prime}(u)[s(s/2+1)u+(1+s)v]\leq 0\quad\text{for all $v\in E$, $s\geq-1$,}$ (10)
@@
$F(u)-F((1+s)u+v)+F^{\prime}(u)[s(s/2+1)u+(1+s)v]\leq 0\qquad\;\text{for all }u,v\in\mathbb{R},s\geq-1$ (11)
@@
$\lambda_{n}<\lambda<\lambda_{n+1},$ (12)
@@
$F_{n}(u)^{+}:=E_{n}\oplus\mathbb{R}^{+}u\>\subset F_{n}(u).$ (14)
@@
$c_{\lambda}:=\inf_{\mathcal{N}\!{\mathcal{P}}_{\lambda}}\mathcal{J}_{\lambda}=\inf_{u\in E\setminus E_{n}}\sup_{F_{n}(u)^{+}}\mathcal{J}_{\lambda}\>>\>0.$ (15)
@@
The key additional abstract result of the present paper is the following theorem, which states that the masses of ground state solutions of (1) form an interval if $\lambda$ is varied in the spectral gap given by (12).
@@
$(0,g_{n})\subseteq M_{n}\subseteq(0,g_{n}]\qquad\text{for some $g_{n}>0$.}$ (16)
We emphazise that Theorem 1.7 yields connectedness of the set $M_{n}$, but it does not yield connectedness of the set of associated ground state solutions. In fact, we do not expect the latter to hold in this general framework.
@@
$g_{n}:=\sup_{\lambda\in(\lambda_{n},\lambda_{n+1})}g(\lambda),$ (17)
@@ 2. Existence of Nehari-Pankov solutions for fixed $\lambda$
$\sup_{v\in E_{n}}\mathcal{J}_{\lambda}(v)\leq\sup_{v\in E_{n}}\frac{q_{\lambda}(v)}{2}\leq 0.\textasciidieresis$ (19)
@@
$0<\sup_{F_{n}(u)^{+}}\mathcal{J}_{\lambda}=\sup_{\overline{F_{n}(u)^{+}}}\mathcal{J}_{\lambda}.$ (20)
@@
$F_{n}(u)^{+}=F_{n}(\tilde{u})^{+}\qquad\text{for every $u\in E\setminus E_{n}$ and every $\tilde{u}\in F_{n}(u)^{+}$.}$ (21)
@@
$\mathcal{J}_{\lambda^{k}}(u_{k})>0\qquad\text{for all $k\in\mathbb{Z}^{>0}$.}$ (22)
@@
$\frac{I(t_{*}w)}{t_{*}^{2}}>\frac{q_{\lambda_{n}}(w)}{2}.$ (23)
@@
$\mathcal{J}_{\lambda}(u)>\mathcal{J}_{\lambda}(w)\qquad\text{for all $w\in F_{n}(u)^{+}\setminus\{u\}$}.$ (24)
@@
$\sup_{k\in\mathbb{Z}^{>0}}\sup_{F_{n}(w_{k})^{+}}\mathcal{J}_{\lambda^{k}}<\infty.$ (25)
@@
$m_{\lambda}(w_{k})\to m_{\lambda}(w)\qquad\text{after passing to a subsequence.}$ (26)
@@
$\mathcal{J}_{\lambda}^{\prime}(v)w<-\varepsilon\qquad\text{for every $v\in B_{\delta}(u)$,}$ (27)
@@
$u_{k}:=m_{\lambda}(w_{k})\to m_{\lambda}(u)=u\qquad\text{as $k\to\infty$.}$ (28)
@@ 3.1. Norms of Nehari-Pankov solutions
$Q_{\lambda}:=\{u\in\mathcal{N}\!{\mathcal{P}}_{\lambda}\mid\mathcal{J}_{\lambda}(u)=c_{\lambda}\},$ (30)
@@
$\limsup_{\varepsilon\to 0^{+}}\frac{c_{\lambda-\varepsilon}-c_{\lambda}}{\varepsilon}\leq\inf\biggl\{\frac{\|u\|_{H}^{2}}{2}\Bigm|u\in Q_{\lambda}\biggr\}$ (31)
@@
$\sup\biggl\{\frac{\|u\|_{H}^{2}}{2}\Bigm|u\in Q_{\lambda}\biggr\}\leq\liminf_{\varepsilon\to 0^{+}}\frac{c_{\lambda}-c_{\lambda+\varepsilon}}{\varepsilon}.$ (32)
@@
$\mathcal{J}_{\lambda-\varepsilon}(w)\leq c_{\lambda}-\kappa\qquad\text{for all $\varepsilon\in(0,\varepsilon_{0})$ and $w\in F^{+}(u)\setminus B_{\delta}(u)$,}$ (33)
@@ 3.2. General action estimates
$\frac{I(t_{n}u_{n})}{t_{n}^{2}}\to 0\qquad\text{as $n\to\infty$.}$ (34)
@@ 4. Further estimates in the case of a metric measure space
$I(u)=\frac{1}{p}\int_{\Omega}r(x)|u|^{p}\,\mathrm{d}x$ (35)
@@
$E$ is compactly embedded in $L^{p}(\Omega)$. (36)
@@
$c_{I}:=\frac{r_{0}\mu(\Omega)^{\frac{2-p}{2}}}{p}.$ (37)
@@
$d(n,\lambda):=\sum_{k=n+1}^{\infty}\frac{1}{\lambda_{k}-\lambda}.$ (38)
@@
$\sup\Bigl\{\|u\|_{L^{2}}\bigm|\text{$u\in Q_{\lambda}$ for some $\lambda\in(\lambda_{n},\lambda_{n+1})$}\Bigr\}\geq\sup_{\lambda\in[\lambda_{n},\lambda_{n+1})}\Bigl[C^{2}C_{p}(\lambda_{n+1}-\lambda)^{\frac{p(p-4)}{(p-2)^{2}}}d(n,\lambda)\Bigr]^{-\frac{p-2}{4}},$ (39)
@@
$C_{p}:=\|r\|_{L^{\infty}(\Omega)}r_{0}^{-\frac{p(p-4)}{(p-2)^{2}}}\mu(\Omega)^{\frac{p-4}{p-2}},$ (40)
@@
$\sup\Bigl\{\|u\|_{L^{2}}\bigm|\text{$u\in Q_{\lambda}$ for some $\lambda\in(\lambda_{n},\lambda_{n+1})$}\Bigr\}\geq\sup_{\lambda\in[\lambda_{n},\lambda_{n+1})}\Bigl[C^{2}C_{p}d(n,\lambda)\Bigr]^{-\frac{1}{p-2}},$ (41)
@@
$C_{p}:=\|r\|_{L^{\infty}(\Omega)}\mu(\Omega)^{\frac{4-p}{2}}.$ (42)
@@
$=C^{2}\Bigl(\sum_{k=n+1}^{\infty}\frac{1}{\lambda_{k}-\lambda}\Bigr)q_{\lambda}(v)=C^{2}d(n,\lambda)q_{\lambda}(v).$ (43)
@@
$c_{\lambda}=\mathcal{J}_{\lambda}(u)=\Bigl(\frac{1}{2}-\frac{1}{p}\Bigr)\int_{\Omega}r(x)|u|^{p}\,\mathrm{d}x\geq\frac{r_{0}(p-2)}{2p}\|u\|_{L^{p}}^{p}.$ (44)
@@
$=C_{p}\|u\|_{L^{2}}^{\frac{4}{p-2}}(\lambda_{n+1}-\lambda)^{\frac{p(p-4)}{(p-2)^{2}}}$ (46)
@@
$\inf_{v\in E_{n}^{\perp}\setminus\{0\}}\frac{1}{\|v\|_{L^{2}}^{2}}\Bigl(q_{\lambda}(v)-\int_{\Omega}V_{u}v^{2}\,\mathrm{d}x\Bigr)\leq 0.$ (47)

### v2

# arXiv:2601.10941v2[math.AP] 01 Mar 2026

@@ 1. Introduction
(1) $Au=\lambda u+I^{\prime}(u),\qquad u\in E$
@@
(2) $\|u\|_{H}=\mu$
for some given $\mu>0$. This constrained problem also has a variational formulation: its solutions are precisely the critical points of the restriction of the energy functional $\mathcal{J}_{0}$ to the sphere $S_{\mu}:=\{u\in H\>:\>\|u\|_{H}=\mu\}$, and the frequency parameter $\lambda$ appears as an a priori unknown Lagrange multiplier.
@@
(3) $u^{\prime\prime}=\lambda u+f(u)$
@@
    (4) $F\Bigl(\frac{s}{\kappa_{0}}\Bigr)\leq F(-s)\leq F(\kappa_{0}s)\qquad\text{for all $s\geq s_{0}$.}$
@@
Very recently, in [16, Theorem 1.3], Jeanjean and Song proved a multiplicity result for compact graphs $\mathcal{G}$ and small masses with continuity-Kirchhoff conditions. More precisely, by combining flow invariance techniques with delicate estimates, they show that for every $n\in\mathbb{Z}^{>0}$ there exists $\mu_{n}>0$ with the property that, for every $\mu\in(0,\mu_{n})$, the NLS equation $u^{\prime\prime}=\lambda u+|u|^{p-2}u$ admits at least $n$ pairs $\pm u$ of sign changing solutions with fixed mass $\mu$. Using our new “frequency to masses” tool (Theorem 1.8 below), we are able to improve this multiplicity result to the existence of infinitely many solutions, for any mass. Moreover, we can extend this multiplicity result beyond the class of odd nonlinearities to a setting where, in particular, Lusternik-Schnirelman theory does not apply. Let us also point out that we can deal with the most general self-adjoint vertex conditions, also generalizing the previous results. In the case of continuity-Kirchhoff conditions, the first eigenfunction of $A$ is positive in $\mathcal{G}$ (see e.g. [18, Theorem 4.12]), so that we also obtain the existence of sign-changing solutions. Finally, as already pointed out, we do not need to make any case distinction depending on whether $p<6$ or $6<p$.
@@
(5) $(-\Delta)^{m}u=\lambda u+f(x,u)\qquad\text{in }\Omega$
@@
(6) $u=\partial_{\nu}u=\dots=\partial_{\nu}^{m-1}u=0\qquad\text{on }\partial\Omega$
@@
(7) $u=\Delta u=\cdots=\Delta^{m-1}u=0\qquad\text{on }\partial\Omega.$
@@
Suppose that (FP1)–(FP4) are satisfied. Then, for any integer $k\geq 1$, there exists $\mu_{k}>0$ such that for all $\mu\in\mathopen{(}0,\mu_{k}\mathclose{)}$, problems (5)–(6) and (5)–(7) both admit $k$ distinct solutions $(u,\lambda)$ with mass $\|u\|_{L^{2}(\Omega)}^{2}=\mu$. Moreover, if $m=1$, these solutions are sign-changing.
We emphasize that this result is new even in the classical second order case $m=1$, and it does not require oddness of the nonlinearity $f$ in $t$. In the special case of homogeneous nonlinearities of the form $f(x,t)=|t|^{p-2}t$, Theorem 1.3 is the analogue of the recent multiplicity result of Jeanjean and Song in [16, Theorem 1.3] for the case of compact metric graphs. Our method is completely different from the approach in [16], which, according to remarks in [16, Introduction], also yields multiplicity results for NLS-equations on bounded domains in $\mathbb{R}^{N}$.
We also point out that the conclusion of Theorem 1.3 is weaker than the one of Theorem 1.1, since only small masses are considered. We conjecture that at least in the case $N=2$, $m>1$ the multiplicity result of Theorem 1.3 extends to arbitrary positive masses $\mu$. To support this conjecture, we present a further result for the closely related problem of finding normalized solutions of biharmonic semilinear equations in the $2$-torus.
@@
(8) $\Delta^{2}u=\lambda u+r(x)|u|^{p-2}u\qquad\text{in $\Omega$}$
@@
Remark 1.5.
(i) In the mass supercritical regime $p>6$ and for a given positive mass $\mu>0$, even the mere existence of one normalized solution of (8) is new. If the weight function $r$ is constant, then the symmetries of the torus can be used to produce multiple solutions with large mass via a reflection procedure (see [21, Theorem 1.13 and Remark 1.1] for related results in bounded domains). This is not possible when $r$ does not have any symmetries.
(ii) Our proof of Theorem 1.4 relies on the explicit knowledge of eigenvalues and eigenfunctions of the Laplacian on the flat torus $S^{1}\times S^{1}$, and it directly extends to the equation
(9) $\Delta^{2}u=\lambda u+r(x)|u|^{p-2}u\qquad\text{in $\Omega=(0,\pi)\times(0,\pi)$}$
subject to Navier boundary conditions
(10) $u=\Delta u=0\qquad\text{on $\partial\Omega$.}$
Here, again we assume that $p>2$ and $r\in L^{\infty}(\Omega)$ satisfies $\underset{x\in\Omega}{\operatorname{essinf}}\,r(x)>0$. Hence, for any $\mu>0$, problem (9), (10) also admits infinitely many normalized sign-changing solutions with mass $\|u\|_{L^{2}(\Omega)}^{2}=\mu$ associated to a sequence of parameters $\lambda$ converging to $\infty$.
@@
    (11) $\lim_{t\to 0^{+}}\frac{I(tu)}{t^{2}}=0\qquad\text{and}\qquad\lim_{t\to+\infty}\frac{I(tu)}{t^{2}}=+\infty.$
@@
    (12) $I(u)-I((1+s)u+v)+I^{\prime}(u)[s(s/2+1)u+(1+s)v]\leq 0\quad\text{for all $v\in E$, $s\geq-1$,}$
@@
(13) $F(u)-F((1+s)u+v)+F^{\prime}(u)[s(s/2+1)u+(1+s)v]\leq 0\qquad\;\text{for all }u,v\in\mathbb{R},s\geq-1$
@@
(14) $\lambda_{n}<\lambda<\lambda_{n+1},$
@@
(16) $F_{n}(u)^{+}:=E_{n}\oplus\mathbb{R}^{+}u\>\subset F_{n}(u).$
@@
(17) $c_{\lambda}:=\inf_{\mathcal{N}\!{\mathcal{P}}_{\lambda}}\mathcal{J}_{\lambda}=\inf_{u\in E\setminus E_{n}}\sup_{F_{n}(u)^{+}}\mathcal{J}_{\lambda}\>>\>0.$
@@
The key additional abstract result of the present paper is the following theorem, which states that the masses of ground state solutions of (1) form a connected set if $\lambda$ is varied in the spectral gap given by (14).
@@
(18) $(0,g_{n})\subseteq M_{n}\subseteq(0,g_{n}]\qquad\text{for some $g_{n}>0$.}$
We emphazise that the connectedness of the set $M_{n}$ stated in Theorem 1.8 does not follow from a standard continuity argument, since we cannot expect connectedness of the set of associated ground state solutions in this general variational framework. Instead, the connectedness of $M_{n}$ follows by an inspection of the special saddle point nature of the Nehari-Pankov sets $\mathcal{N}\!{\mathcal{P}}_{\lambda}$, while it is not a sole consequence of the continuity of ground state levels $c_{\lambda}$.
@@
(19) $g_{n}:=\sup_{\lambda\in(\lambda_{n},\lambda_{n+1})}g(\lambda),$
@@ 2. Existence of Nehari-Pankov solutions for fixed $\lambda$
(21) $\sup_{v\in E_{n}}\mathcal{J}_{\lambda}(v)\leq\sup_{v\in E_{n}}\frac{q_{\lambda}(v)}{2}\leq 0.\textasciidieresis$
@@
(22) $0<\sup_{F_{n}(u)^{+}}\mathcal{J}_{\lambda}=\sup_{\overline{F_{n}(u)^{+}}}\mathcal{J}_{\lambda}.$
@@
(23) $F_{n}(u)^{+}=F_{n}(\tilde{u})^{+}\qquad\text{for every $u\in E\setminus E_{n}$ and every $\tilde{u}\in F_{n}(u)^{+}$.}$
@@
(24) $\mathcal{J}_{\lambda^{k}}(u_{k})>0\qquad\text{for all $k\in\mathbb{Z}^{>0}$.}$
@@
(25) $\frac{I(t_{*}w)}{t_{*}^{2}}>\frac{q_{\lambda_{n}}(w)}{2}.$
@@
(26) $\mathcal{J}_{\lambda}(u)>\mathcal{J}_{\lambda}(w)\qquad\text{for all $w\in F_{n}(u)^{+}\setminus\{u\}$}.$
@@
(27) $\sup_{k\in\mathbb{Z}^{>0}}\sup_{F_{n}(w_{k})^{+}}\mathcal{J}_{\lambda^{k}}<\infty.$
@@
(28) $m_{\lambda}(w_{k})\to m_{\lambda}(w)\qquad\text{after passing to a subsequence.}$
@@
(29) $\mathcal{J}_{\lambda}^{\prime}(v)w<-\varepsilon\qquad\text{for every $v\in B_{\delta}(u)$,}$
@@
(30) $u_{k}:=m_{\lambda}(w_{k})\to m_{\lambda}(u)=u\qquad\text{as $k\to\infty$.}$
@@ 3.1. Norms of Nehari-Pankov solutions
(32) $Q_{\lambda}:=\{u\in\mathcal{N}\!{\mathcal{P}}_{\lambda}\mid\mathcal{J}_{\lambda}(u)=c_{\lambda}\},$
@@
(33) $\limsup_{\varepsilon\to 0^{+}}\frac{c_{\lambda-\varepsilon}-c_{\lambda}}{\varepsilon}\leq\inf\biggl\{\frac{\|u\|_{H}^{2}}{2}\Bigm|u\in Q_{\lambda}\biggr\}$
@@
(34) $\sup\biggl\{\frac{\|u\|_{H}^{2}}{2}\Bigm|u\in Q_{\lambda}\biggr\}\leq\liminf_{\varepsilon\to 0^{+}}\frac{c_{\lambda}-c_{\lambda+\varepsilon}}{\varepsilon}.$
@@
(35) $\mathcal{J}_{\lambda-\varepsilon}(w)\leq c_{\lambda}-\kappa\qquad\text{for all $\varepsilon\in(0,\varepsilon_{0})$ and $w\in F^{+}(u)\setminus B_{\delta}(u)$,}$
@@ 3.2. General action estimates
(36) $\frac{I(t_{n}u_{n})}{t_{n}^{2}}\to 0\qquad\text{as $n\to\infty$.}$
@@ 4. Further estimates in the case of a metric measure space
(37) $I(u)=\frac{1}{p}\int_{\Omega}r(x)|u|^{p}\,\mathrm{d}x$
@@
(38) $E$ is compactly embedded in $L^{p}(\Omega)$.
@@
(39) $c_{I}:=\frac{r_{0}\mu(\Omega)^{\frac{2-p}{2}}}{p}.$
@@
(40) $d(n,\lambda):=\sum_{k=n+1}^{\infty}\frac{1}{\lambda_{k}-\lambda}.$
@@
(41) $\sup\Bigl\{\|u\|_{L^{2}}\bigm|\text{$u\in Q_{\lambda}$ for some $\lambda\in(\lambda_{n},\lambda_{n+1})$}\Bigr\}\geq\sup_{\lambda\in[\lambda_{n},\lambda_{n+1})}\Bigl[C^{2}C_{p}(\lambda_{n+1}-\lambda)^{\frac{p(p-4)}{(p-2)^{2}}}d(n,\lambda)\Bigr]^{-\frac{p-2}{4}},$
@@
(42) $C_{p}:=\|r\|_{L^{\infty}(\Omega)}r_{0}^{-\frac{p(p-4)}{(p-2)^{2}}}\mu(\Omega)^{\frac{p-4}{p-2}},$
@@
(43) $\sup\Bigl\{\|u\|_{L^{2}}\bigm|\text{$u\in Q_{\lambda}$ for some $\lambda\in(\lambda_{n},\lambda_{n+1})$}\Bigr\}\geq\sup_{\lambda\in[\lambda_{n},\lambda_{n+1})}\Bigl[C^{2}C_{p}d(n,\lambda)\Bigr]^{-\frac{1}{p-2}},$
@@
(44) $C_{p}:=\|r\|_{L^{\infty}(\Omega)}\mu(\Omega)^{\frac{4-p}{2}}.$
@@
(45) $=C^{2}\Bigl(\sum_{k=n+1}^{\infty}\frac{1}{\lambda_{k}-\lambda}\Bigr)q_{\lambda}(v)=C^{2}d(n,\lambda)q_{\lambda}(v).$
@@
(46) $c_{\lambda}=\mathcal{J}_{\lambda}(u)=\Bigl(\frac{1}{2}-\frac{1}{p}\Bigr)\int_{\Omega}r(x)|u|^{p}\,\mathrm{d}x\geq\frac{r_{0}(p-2)}{2p}\|u\|_{L^{p}}^{p}.$
@@
(48) $=C_{p}\|u\|_{L^{2}}^{\frac{4}{p-2}}(\lambda_{n+1}-\lambda)^{\frac{p(p-4)}{(p-2)^{2}}}$
@@
(49) $\inf_{v\in E_{n}^{\perp}\setminus\{0\}}\frac{1}{\|v\|_{L^{2}}^{2}}\Bigl(q_{\lambda}(v)-\int_{\Omega}V_{u}v^{2}\,\mathrm{d}x\Bigr)\leq 0.$